# A Linear Bound on the Diameter of the Kakimizu Complex for Hyperbolic Knots

### v3

@@ 1 Introduction
Given a knot $K$ (or a link $L$) in $\mathbb{S}^{3}$, a spanning surface is an embedded (possibly disconnected) compact surface whose boundary is exactly $K$ (or $L$). A connected, orientable spanning surface is called a Seifert surface. A Seifert surface of minimal genus is called a minimal-genus Seifert surface, and its genus is referred to as the genus of the knot (or link), denoted by $g(K)$ (often abbreviated to $g$ in this paper). The Kakimizu complex $MS(K)$ of a knot $K\subset\mathbb{S}^{3}$, introduced by Kakimizu in [Kak92], encodes both the classification of minimal-genus Seifert surfaces for $K$ and the topological adjacency relations among them.
The topology and simplicial structure of $MS(K)$ have been extensively studied: Kakimizu showed that it is connected [Kak92], Schultens proved that it is simply connected [Sch10], and Przytycki–Schultens later established its contractibility [PS12]. Moreover, Agol–Zhang showed that every saturated simplex of $MS(K)$ has maximal dimension [AZ22]. In certain families of knots, the complex can even be computed explicitly [Kak92, Kak05, Sak94].
In the study of the Kakimizu complex diameter, Pelayo [Pel07] and Sakuma–Shackleton [SS09], using different methods, established a quadratic upper bound for the diameter of the Kakimizu complex of an atoroidal knot in terms of the genus of the knot. In this paper, we improve this result by proving a linear upper bound for the diameter of the Kakimizu complex of an atoroidal knot, thereby resolving a conjecture posed by Sakuma–Shackleton [SS09, Question 1.7]:
@@
The example from [SS09, Proposition 1.4] shows that the bound is sharp when $g=1$, where the diameter equals $2$. Moreover, the bound is optimal in order, as there exist atoroidal knots whose Kakimizu complexes have diameter at least 2g12g-1 [SS09, Proposition 1.6].
According to Thurston’s trichotomy theorem [Mar16, Proposition 11.2.11], atoroidal knots include both torus knots and hyperbolic knots. The remaining knots, those that are not atoroidal, are referred to as satellite knots. A torus knot has a unique incompressible Seifert surface [Tsa94], so its Kakimizu complex is trivial. Therefore, Theorem˜1.1 primarily focuses on hyperbolic knots. Moreover, for hyperbolic knots, every vertex in the Kakimizu complex has finite valence [Ban11, Theorem 1.8]. Combining these facts, hyperbolic knots admit only finitely many minimal-genus Seifert surfaces. In contrast, the properties mentioned above do not necessarily hold for satellite knots: the isotopy classes of their Seifert surfaces may form an infinite set [Eis77], the Kakimizu complex may have infinite diameter [Kak92], and some vertices may have infinite valence [Ban11].
Beyond the complex $MS(K)$ consisting of minimal-genus Seifert surfaces, Kakimizu [Kak92] also introduced a complex $IS(K)$ generated by incompressible Seifert surfaces. Although every minimal-genus Seifert surface is incompressible, the complexes $MS(K)$ and $IS(K)$ do not always coincide. For instance, certain pretzel knots are hyperbolic and admit infinitely many incompressible Seifert surfaces of unbounded genus [Wil08], whereas a hyperbolic knot admits only finitely many minimal-genus Seifert surfaces.
@@
Our approach is inspired by the work of Sakuma–Shackleton [SS09], who analysed the structure of lifts of Seifert surfaces in the infinite cyclic cover of $\mathbb{S}^{3}\setminus K$. Concretely, suppose that an atoroidal knot $K$ admits two distinct minimal-genus Seifert surfaces $S_{1}$ and $S_{2}$. Fix a lift of $S_{1}$ to the infinite cyclic cover. Then the lifts of $S_{2}$ decompose this lift of $S_{1}$ into a linearly ordered sequence of subsurfaces, and Sakuma–Shackleton established the following properties:
@@
While the initial observations yield only a quadratic upper bound for $\operatorname{diam}(MS(K))$, the key new ingredient of this paper is a refinement of this analysis via annulus surgery. Using this technique, we show that all subsurfaces with negative Euler characteristic are contained in a sequence of consecutive subsurfaces of linear size in the genus, which we call the central zone. (see Section˜7). This additional structure allows us to control the configuration much more effectively and reduces the resulting diameter bound to linear order.
@@ 2.1 Kakimizu’s complexes
Definition 2.1 (Kakimizu complex, [Kak92]).
@@
Definition 2.2 (Incompressible Seifert surface complex, [Kak92]).
@@
Definition 2.5 (Natural distance, [SS09]).
@@ 2.2 Atoroidal knots and hyperbolic knots
Thurston’s trichotomy theorem states that every atoroidal knot is either a torus knot or a hyperbolic knot [Mar16, Proposition 11.2.11]. Therefore, every knot is either a torus knot, a hyperbolic knot, or a satellite knot.
@@
For a torus knot $K$, since it is fibered, it admits a unique minimal-genus Seifert surface given by the fiber. Consequently, its Kakimizu complex $MS(K)$ is trivial. Moreover, by [Tsa94], a torus knot admits a unique incompressible Seifert surface up to isotopy, and hence both $IS_{\ell}(K)$ and $IS(K)$ are trivial. Therefore, when considering upper bounds on the diameter of $MS(K)$, $IS_{\ell}(K)$, and $IS(K)$ for atoroidal knots, it suffices to restrict attention to the hyperbolic case.
@@ 3 Intersection of Two Seifert Surfaces
If $\gamma$ is boundary-parallel on $S_{2}$, since $S_{1}$ is minimal-genus, by ˜3.3, $\gamma$ must also be boundary-parallel on $S_{1}$. We may suppose that $\gamma$ and $K$ border two annuli $A_{1}\subset S_{1}$ and $A_{2}\subset S_{2}$. The union $A_{1}\cup A_{2}$, denoted by $\mathbb{T}$, is an embedded torus in $\mathbb{S}^{3}$ such that $K\subset\mathbb{T}$. Cutting $\mathbb{S}^{3}$ along $\mathbb{T}$ results in two components, $X_{1}$ and $X_{2}$. In what follows, we assume (without loss of generality) that $X_{1}$ is a solid torus (see, e.g., [Hat07, Section 1.1, Exercise 2]).
@@
As an immediate consequence of this lemma, we obtain the following corollary. In fact, this result already appears as [SS09, Lemma 3.2].
Corollary 3.5 ([SS09, Lemma 3.2]).
@@
However, it is still unknown whether two incompressible but non-minimal-genus Seifert surfaces for an atoroidal knot can be isotoped so that every intersection loop is essential on both surfaces (see [OT03, Question 1.6]). If this were true, the bound in Theorem˜1.2 could be improved by a factor of two.
@@ 4 A Technical Lemma
Let $K$ be an atoroidal knot, and let $(S_{1},S_{2})$ be a favourable pair. Suppose that there exist two connected components $C_{1}$ and $C_{2}$ of $S_{1}\cap S_{2}$, both of which are simple loops, such that $C_{1}$ and $C_{2}$ are isotopic on $S_{1}$ as well as on $S_{2}$. For $i=1,2$, let $A_{i}\subset S_{i}$ be the annulus bounded by $C_{1}$ and $C_{2}$. We define $\mathbb{T}:=A_{1}\cup A_{2}$, which is an embedded torus in $\mathbb{S}^{3}\setminus K$. Then $\mathbb{T}$ cannot be boundary-parallel in $\mathbb{S}^{3}\setminus K$.
@@
Because the $1$-manifold $R\cap\mathbb{T}$ lies on the annulus $A_{1}$, the homology class $i_{*}([R\cap A_{1}])$ must be a multiple of $C_{1}]=[C_{2}]$ in $H_{1}(X_{1})$. Combined with the fact that $i_{*}([R\cap\mathbb{T}])$ generates $H_{1}(X_{1})$, this implies that both $C_{1}]$ and $C_{2}]$ generate $H_{1}(X_{1})$. Consequently, the pair $(X_{1},A_{1})$ admits a product structure $(X_{1},A_{1})\cong\left(A_{1}\times[0,1],\,A_{1}\times\{0\}\right)$.
@@
Moreover, since $R$ cannot be a disk (otherwise $R$ and $A_{2}$ would lie in different connected components of $S_{2}$, contradicting the connectedness of $S_{2}$), we conclude that $R$ is an annulus. This implies that $K$ and $C_{3}:=R\cap A_{1}\subset S_{1}\cap S_{2}$ are isotopic on $S_{2}$, which contradicts the fact that every intersection loop on $S_{2}$ is not boundary-parallel (except for $K$). ∎
@@ 5.1 Definition of infinite cyclic cover
Let $K\subset\mathbb{S}^{3}$ be a knot with tubular neighbourhood $W(K)$. The infinite cyclic cover $\widetilde{E}$ of the knot exterior $E=\mathbb{S}^{3}\setminus W(K)$ is the covering space associated to the kernel of the abelianisation $\mathrm{ab}:\pi_{1}(E)\rightarrow H_{1}(E)\cong\mathbb{Z}$.
@@
Let $S$ be a Seifert surface for $K$. When lifted to the infinite cyclic cover $\widetilde{E}$, we pick one lift as $(\widetilde{S})_{0}$, and for any integer $i\in\mathbb{Z}$, define $(\widetilde{S})_{i}=\tau^{i}((\widetilde{S})_{0})$, where $\tau$ is the generator of the deck transformation group. So $(\widetilde{S})_{i}:i\in\mathbb{Z}$ are all lifts of $S$.
The collection of surfaces $\{(\widetilde{S})_{i}\}_{i\in\mathbb{Z}}$ divides $\widetilde{E}$ into countably many regions $\{\widetilde{E}_{i}(S)\}_{i\in\mathbb{Z}}$, where each $\widetilde{E}_{i}(S)$ is divided by $(\widetilde{S})_{i}$ and $(\widetilde{S})_{i+1}$. (See Fig.˜2 for illustration.)
@@ 5.2 Infinite cyclic cover decomposition
Definition 5.9.
Let $(\widetilde{S_{1}})_{i}^{j}\subset(\widetilde{S_{1}})_{i}$ be defined as $(\widetilde{S_{1}})_{i}\cap\widetilde{E}_{j}(S_{2})$, and similarly, let $(\widetilde{S_{2}})_{j}^{i}\subset(\widetilde{S_{2}})_{j}$ be defined as $(\widetilde{S_{2}})_{j}\cap\widetilde{E}_{i}(S_{1})$.
By definition we have
@@ 5.3 Kakimizu’s characterisation
Definition 5.12 (Kakimizu’s characterisation, [SS09]).
@@
    ([Kak92, Proposition 3.1]) For any two vertices $S_{1}]$ and $S_{2}]$ with both genus at most $\ell$, we have
@@
Kakimizu [Kak92] proved the first assertion of this theorem, while the proof of the second assertion essentially follows Kakimizu’s original argument [Kak92]. It is straightforward to extend this argument to the setting of $IS_{\ell}(K)$ provided that $S_{1}$ has minimal genus. Indeed, Kakimizu constructed a path in $IS_{\infty}(K)$ connecting $S_{1}]$ and $S_{2}]$. Since $S_{1}$ has minimal genus, every Seifert surface appearing along this path has genus at most $\ell$, and hence the same path lies entirely in $IS_{\ell}(K)$. For completeness, we present this proof below to make this point explicit.
@@
By Theorem˜5.14, we immediately obtain the following: for any two vertices $S_{1}]$ and $S_{2}]$ in $IS(K)$ of minimal genus $g=g(K)$, we have $d_{g}([S_{1}],[S_{2}])=d_{\infty}([S_{1}],[S_{2}])=d_{*}([S_{1}],[S_{2}])$. In fact, this result was already proved by Kakimizu in [Kak92, Proposition 3.1].

### v4

@@ 1 Introduction
Given a knot $K$ (or a link $L$) in $\mathbb{S}^{3}$, a spanning surface is an embedded (possibly disconnected) compact surface whose boundary is exactly $K$ (or $L$). A connected, orientable spanning surface is called a Seifert surface. A Seifert surface of minimal genus is called a minimal-genus Seifert surface, and its genus is referred to as the genus of the knot (or link), denoted by $g(K)$ (often abbreviated to $g$ in this paper). The Kakimizu complex $MS(K)$ of a knot $K\subset\mathbb{S}^{3}$, introduced by Kakimizu in [6], encodes both the classification of minimal-genus Seifert surfaces for $K$ and the topological adjacency relations among them.
The topology and simplicial structure of $MS(K)$ have been extensively studied: Kakimizu showed that it is connected [6], Schultens proved that it is simply connected [13], and Przytycki–Schultens later established its contractibility [10]. Moreover, Agol–Zhang showed that every saturated simplex of $MS(K)$ has maximal dimension [1]. In certain families of knots, the complex can even be computed explicitly [6, 7, 12].
In the study of the Kakimizu complex diameter, Pelayo [9] and Sakuma–Shackleton [11], using different methods, established a quadratic upper bound for the diameter of the Kakimizu complex of an atoroidal knot in terms of the genus of the knot. In this paper, we improve this result by proving a linear upper bound for the diameter of the Kakimizu complex of an atoroidal knot, thereby resolving a conjecture posed by Sakuma–Shackleton [11, Question 1.7]:
@@
The example from [11, Proposition 1.4] shows that the bound is sharp when $g=1$, where the diameter equals $2$. Moreover, the bound is optimal in order, as there exist atoroidal knots whose Kakimizu complexes have diameter at least 2g12g-1 [11, Proposition 1.6].
According to Thurston’s trichotomy theorem [8, Proposition 11.2.11], atoroidal knots include both torus knots and hyperbolic knots. The remaining knots, those that are not atoroidal, are referred to as satellite knots. A torus knot has a unique incompressible Seifert surface [14], so its Kakimizu complex is trivial. Therefore, Theorem˜1.1 primarily focuses on hyperbolic knots. Moreover, for hyperbolic knots, every vertex in the Kakimizu complex has finite valence [2, Theorem 1.8]. Combining these facts, hyperbolic knots admit only finitely many minimal-genus Seifert surfaces (however, the number of minimal-genus Seifert surfaces of a hyperbolic knot cannot be bounded solely as a function of the genus, as shown by Tsutsumi [16, Theorem 1.1]).
In contrast, the properties mentioned above do not necessarily hold for satellite knots: the isotopy classes of their Seifert surfaces may form an infinite set [3], the Kakimizu complex may have infinite diameter [6], and some vertices may have infinite valence [2].
Beyond the complex $MS(K)$ consisting of minimal-genus Seifert surfaces, Kakimizu [6] also introduced a complex $IS(K)$ generated by incompressible Seifert surfaces. Although every minimal-genus Seifert surface is incompressible, the complexes $MS(K)$ and $IS(K)$ do not always coincide. For instance, certain pretzel knots are hyperbolic and admit infinitely many incompressible Seifert surfaces of unbounded genus [17], whereas a hyperbolic knot admits only finitely many minimal-genus Seifert surfaces.
@@
Our approach is inspired by the work of Sakuma–Shackleton [11], who analysed the structure of lifts of Seifert surfaces in the infinite cyclic cover of $\mathbb{S}^{3}\setminus K$. Concretely, suppose that an atoroidal knot $K$ admits two distinct minimal-genus Seifert surfaces $S_{1}$ and $S_{2}$. Fix a lift of $S_{1}$ to the infinite cyclic cover. Then the lifts of $S_{2}$ decompose this lift of $S_{1}$ into a linearly ordered sequence of subsurfaces, and Sakuma–Shackleton established the following properties:
@@
While the initial observations yield only a quadratic upper bound for $\operatorname{diam}(MS(K))$, the key new ingredient of this paper is a refinement of this analysis via annulus surgery. Using this technique, we show that all subsurfaces with negative Euler characteristic are contained in a sequence of consecutive subsurfaces of linear size in the genus, which we call the central zone (see Section˜7). This additional structure allows us to control the configuration much more effectively and reduces the resulting diameter bound to linear order.
@@ 2.1 Kakimizu’s complexes
Definition 2.1 (Kakimizu complex, [6]).
@@
Definition 2.2 (Incompressible Seifert surface complex, [6]).
@@
Definition 2.5 (Natural distance, [11]).
@@ 2.2 Atoroidal knots and hyperbolic knots
Thurston’s trichotomy theorem states that every atoroidal knot is either a torus knot or a hyperbolic knot [8, Proposition 11.2.11]. Therefore, every knot is either a torus knot, a hyperbolic knot, or a satellite knot.
@@
For a torus knot $K$, since it is fibered, it admits a unique minimal-genus Seifert surface given by the fiber. Consequently, its Kakimizu complex $MS(K)$ is trivial. Moreover, by [14], a torus knot admits a unique incompressible Seifert surface up to isotopy, and hence both $IS_{\ell}(K)$ and $IS(K)$ are trivial. Therefore, when considering upper bounds on the diameter of $MS(K)$, $IS_{\ell}(K)$, and $IS(K)$ for atoroidal knots, it suffices to restrict attention to the hyperbolic case.
@@ 3 Intersection of Two Seifert Surfaces
If $\gamma$ is boundary-parallel on $S_{2}$, without loss of generality, we may assume that $\gamma$ is the boundary-parallel loop on $S_{2}$ that is closest to the knot $K$. Since $S_{1}$ has minimal genus, by ˜3.3, the loop $\gamma$ must also be boundary-parallel on $S_{1}$. We may suppose that $\gamma$ and $K$ cobound two annuli $A_{1}\subset S_{1}$ and $A_{2}\subset S_{2}$. The union $A_{1}\cup A_{2}$, denoted by $\mathbb{T}$, is a torus in $\mathbb{S}^{3}$ such that $K\subset\mathbb{T}$. Moreover, since $\gamma$ is chosen to be the boundary-parallel loop on $S_{2}$ closest to the knot $K$, the annulus $A_{2}$ intersects $A_{1}$ only along $K$ and $\gamma$. Hence, $\mathbb{T}$ is an embedded torus. Cutting $\mathbb{S}^{3}$ along $\mathbb{T}$ results in two components, $X_{1}$ and $X_{2}$. In what follows, we assume (without loss of generality) that $X_{1}$ is a solid torus (see, e.g., [5, Section 1.1, Exercise 2]).
@@
As an immediate consequence of this lemma, we obtain the following corollary. In fact, this result already appears as [11, Lemma 3.2].
Corollary 3.5 ([11, Lemma 3.2]).
@@
Remark 3.6.
In the definition of a favourable pair, we require that except for $K$, every intersection loop is essential on $S_{2}$, in particular, not boundary-parallel on $S_{2}$. This requirement is necessary because this property is used essentially in Lemma˜4.1 and Proposition˜8.1. Moreover, since it suffices for only one of the two surfaces to satisfy this condition, the definition of a favourable pair is chosen to be just strong enough for our purposes.
To ensure that this property holds up to isotopy, we must require that $S_{1}$ has minimal genus; otherwise, we cannot always arrange for $S_{2}$ to satisfy the above condition. Indeed, [15, Theorem 1.1] provides infinitely many examples of incompressible but non-minimal-genus Seifert surfaces for a hyperbolic knot which contain a loop $\gamma$ that is not boundary-parallel on $S_{1}$ but is boundary-parallel in $\mathbb{S}^{3}\setminus K$. Such surfaces are referred to as accidental. Consequently, if $S_{1}$ does not have minimal genus and is such an accidental surface, and if $S_{2}$ intersects $S_{1}$ along the loop $\gamma$, then, provided that $S_{2}$ is not accidental, this loop may become parallel to $K$ on $S_{2}$.
@@ 4 A Technical Lemma
Let $K$ be an atoroidal knot, and let $(S_{1},S_{2})$ be a favourable pair. Suppose that there exist two connected components $C_{1}$ and $C_{2}$ of $S_{1}\cap S_{2}$, both of which are simple loops, such that $C_{1}$ and $C_{2}$ are isotopic on $S_{1}$ as well as on $S_{2}$. For $i=1,2$, let $A_{i}\subset S_{i}$ be the annulus bounded by $C_{1}$ and $C_{2}$. We define $\mathbb{T}:=A_{1}\cup A_{2}$, which is a torus in $\mathbb{S}^{3}\setminus K$. If $\mathbb{T}$ is embedded, then $\mathbb{T}$ cannot be boundary-parallel in $\mathbb{S}^{3}\setminus K$.
@@
Because the $1$-manifold $R\cap\mathbb{T}$ lies on the annulus $A_{1}$ and $C_{1}$, $C_{2}$ cobound $A_{1}$, the homology class $i_{*}([R\cap A_{1}])$ must be a multiple of $C_{1}]=[C_{2}]$ in $H_{1}(X_{1})$. Combined with the fact that $i_{*}([R\cap\mathbb{T}])$ generates $H_{1}(X_{1})$, this implies that both $C_{1}]$ and $C_{2}]$ are equal to $i_{*}([R\cap A_{1}])$ and they generate $H_{1}(X_{1})$. Consequently, the pair $(X_{1},A_{1})$ admits a product structure $(X_{1},A_{1})\cong\left(A_{1}\times[0,1],\,A_{1}\times\{0\}\right)$.
@@
Moreover, since $R$ cannot be a disk (otherwise $R$ and $A_{2}$ would lie in different connected components of $S_{2}$, contradicting the connectedness of $S_{2}$), we conclude that $R$ is an annulus. This implies that $K$ and $C_{3}:=R\cap A_{1}\subset S_{1}\cap S_{2}$ are isotopic on $S_{2}$, which contradicts the fact that except for $K$, every intersection loop on $S_{2}$ is not boundary-parallel (see Definition˜3.2). ∎
@@ 5.1 Definition of infinite cyclic cover
Let $K\subset\mathbb{S}^{3}$ be a knot with open tubular neighbourhood $W(K)$. The infinite cyclic cover $\widetilde{E}$ of the knot exterior $E=\mathbb{S}^{3}\setminus W(K)$ is the covering space associated to the kernel of the abelianisation $\mathrm{ab}:\pi_{1}(E)\rightarrow H_{1}(E)\cong\mathbb{Z}$.
@@
Let $S$ be a Seifert surface for $K$. From now on, when no confusion arises, we do not distinguish between $S$ and $S\setminus W(K)$, and we denote both simply by $S$. When lifted to the infinite cyclic cover $\widetilde{E}$, we pick one lift as $(\widetilde{S})_{0}$, and for any integer $i\in\mathbb{Z}$, define $(\widetilde{S})_{i}=\tau^{i}((\widetilde{S})_{0})$, where $\tau$ is the generator of the deck transformation group. So $(\widetilde{S})_{i}:i\in\mathbb{Z}$ are all lifts of $S$.
The collection of surfaces $\{(\widetilde{S})_{i}\}_{i\in\mathbb{Z}}$ divides $\widetilde{E}$ into countably many $3$-dimensional regions $\{\widetilde{E}_{i}(S)\}_{i\in\mathbb{Z}}$, where each $\widetilde{E}_{i}(S)$ is divided by $(\widetilde{S})_{i}$ and $(\widetilde{S})_{i+1}$. (See Fig.˜2 for illustration.)
@@ 5.2 Infinite cyclic cover decomposition
Definition 5.5.
Let $(\widetilde{S_{1}})_{i}^{j}\subset(\widetilde{S_{1}})_{i}$ be defined as $(\widetilde{S_{1}})_{i}\cap\widetilde{E}_{j}(S_{2})$, which is a subsurface of the surface $(\widetilde{S_{1}})_{i}$, and similarly, let $(\widetilde{S_{2}})_{j}^{i}\subset(\widetilde{S_{2}})_{j}$ be defined as $(\widetilde{S_{2}})_{j}\cap\widetilde{E}_{i}(S_{1})$.
@@ 5.3 Kakimizu’s characterisation
Definition 5.12 (Kakimizu’s characterisation, [11]).
@@
    ([6, Proposition 3.1]) For any two vertices $S_{1}]$ and $S_{2}]$ with both genus at most $\ell$, we have
@@
Kakimizu [6] proved the first assertion of this theorem, while the proof of the second assertion essentially follows Kakimizu’s original argument [6]. It is straightforward to extend this argument to the setting of $IS_{\ell}(K)$ provided that $S_{1}$ has minimal genus. Indeed, Kakimizu constructed a path in $IS_{\infty}(K)$ connecting $S_{1}]$ and $S_{2}]$. Since $S_{1}$ has minimal genus, every Seifert surface appearing along this path has genus at most $\ell$, and hence the same path lies entirely in $IS_{\ell}(K)$. For completeness, we present this proof below to make this point explicit.
@@
By Theorem˜5.14, we immediately obtain the following: for any two vertices $S_{1}]$ and $S_{2}]$ in $IS(K)$ of minimal genus $g=g(K)$, we have $d_{g}([S_{1}],[S_{2}])=d_{\infty}([S_{1}],[S_{2}])=d_{*}([S_{1}],[S_{2}])$. In fact, this result was already proved by Kakimizu in [6, Proposition 3.1].